INTERPRETATION OF THE IR/UV SPECTRA OF Ac-Trp-Tyr-NH_{2} and Ac-Trp-Tyr-Ser-NH_{2} USING MOLECULAR DYNAMICS AND AB INITIO METHODS.^{a}

\underline{JESSICA A. THOMAS} and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260; ERIC GLOAGUEN, BENJAMIN TARDIVEL, FRANÇOIS PIUZZI, and MICHEL MONS, Laboiratoire Francis Perrin, URA 2453 CRNS, Service des Photons, Atomes et Molécules CEA Saclay, Bât 522, 91191 Gif-sur-Yvette Cedex, France..

The peptides Ac-Trp-Tyr-NH_{2} and Ac-Trp-Tyr-Ser-NH_{2}, which form the N-terminal region of a folding nucleus in βlactoglobulin, were studied in the gas phase using IR/UV double resonance spectroscopy and initial results were presented at a previous symposium.^{b} Molecular dynamics (AMBER 99/99SB, CHARMM 27) and ab initio calculations (RI-B97-D/TZVPP, pbe GGA/cc-PVDZ) resulted in an improved interpretation of the spectra and assignments for the observed conformers. Results are compared to similar molecules such as Ac-Trp-NH_{2} and Ac-Phe-Phe-NH_{2}.

aWork supported in part by NSF CHE-0911117

bJ.A. Thomas, D.W. Pratt, E. Gloaguen, B. Tardivel, F. Piuzzi, and M. Mons 63^{rd} International Symposium on Molecular Spectroscopy, FB11, 2008.